# Building Symbiotic Artificial Intelligence: Reviewing the AI Act for a Human-Centred, Principle-Based Framework

### v2

@@ 1 Introduction
This work aims to identify the principles that can guide the design of SAI systems to comply with the current legal landscape considering humans as the core of the development process [29]. With this goal in mind, we performed a Systematic Literature Review (SLR), employing content analysis to understand what researchers suggest about the relationship between SAI and the AI Act. The main contribution of this study is the proposal of principles to guide stakeholders in the design of SAI systems, i.e., of HCAI systems that favour a symbiotic relationship with humans. In addition, three properties strictly connected to the principles emerged that are represented in a value-based framework shown in Figure 3. It is finally put in the context of the relevant literature, comparing it with other principled-based approaches to SAI and HCAI. The goal of this study is thus better formulated in the following research question:
@@
The article is structured as follows: Section 2 describes the method adopted to conduct the SLR; Section 3 presents the results obtained by the SLR (i.e., principles and properties that characterize the human-AI symbiosis); Section 4 discusses the analysis of the literature presenting the challenges, trends, and limitations faced while conducting the SLR.
2 Methods
@@
    european AND (“artificial intelligence” OR ai) AND act AND (human OR human-centric OR human-centric OR human-centric)
@@
The latter was used to build the queries to run on Scopus because it is comprehensive and includes the most relevant and accredited digital libraries, including journal and conference manuscripts.
@@ 2.2 Coding and Classification
The papers were classified using a mixed approach of data coding: i.e., a-priori and in-vivo. A-priori coding consists of the categorisation of papers with dimensions that are established before the classification process. In contrast, in-vivo coding involves the definition of the dimensions as the papers are read and analysed [36]. The name of each principle was collectively chosen by the researchers who took part in the SLR, reflecting terms that are recurrent in the set of papers and those used in the regulation.
@@ 2.3 Conducting the SLR
The first step of the execution was searching for academic articles about the topic by running the five queries on the Scopus digital library. Each query provided the following results: Q1: 530, Q2: 223, Q3: 4, Q4: 1, and Q5: 4. It emerged that Q1 and Q2 returned a higher number of papers obtained since they are broader and more exploratory than the others.
@@
From a total of 762 papers, after a check for duplicates (i.e., 158 files), 604 publications were obtained. A more in-depth step was taken, which resulted in the removal of 365 papers since they did not match the inclusion criteria. In the end, each publication was further analysed by reviewing the abstract, the introduction, and the conclusions, obtaining the final set of papers. The final set used for the literature review contains 58 papers. Figure 1 shows the PRISMA diagram of the whole process [56].
@@
3 Results: A Principled Human-Centred AI Framework
Through content analysis performed on 58 articles, four principles were identified as requirements for designing AI Act-compliant HCAI systems: Transparency, Fairness, Automation Level, and Protection. These are described in the following sections and are represented in Figure 2. Through further analysis, three additional properties were identified: Trustworthiness, Robustness, and Sustainability.
The four principles and their relationship with the additional three properties are detailed in the following sections, discussing how they can support the creation of AI systems, fostering a symbiotic relationship with humans [11].
@@ 3.1 Principles for Human-AI Symbiosis
Defining principles to guide the design and development of AI systems through a human-centred approach is crucial to establish a symbiotic relationship between the two parties [11]. It is important to guarantee that AI systems adhere to the highest standards of ethical conduct. Thus, designers and developers must ensure that the implemented solutions comply with laws and regulations safeguarding human rights and European values, making these systems safer for end-users [21, 79]. To this end, providers must register high-risk AI models in an EU database managed by the European Commission to enhance public transparency and enable oversight by authorities [58].
@@ 3.1.1 Transparency
Transparency can guarantee that AI systems are effectively overseen by humans and allow intervention when potential harm occurs [58]. Transparency ensures that critical information about how the AI model was trained and structured is available to humans [58], since stakeholders must understand how AI models function and the reasoning behind their decisions to be able to intervene [58]. Being a multi-faceted property that concerns AI models, their components, and algorithms, Transparency serves as an umbrella term to the functional understanding of the model and the rationale behind its operations [18, 75]. It is characterized by two dimensions: Explainability and Interpretability [24].
@@
Explainability aims to provide explanations concerning AI systems’ operations, ensuring that even when humans cannot understand how an AI system provides an output, they can at least receive information about why it was produced [67]. In other words, this aspect contributes to the degree to which an AI system is open and observable to humans [58, 75]. When interacting with AI systems, humans should have the right to explanations to make informed decisions, supported by AI and not replaced by it [39, 78], while being able to oversee its processes [74].
@@
An AI system can be considered interpretable if it can be correctly understood by an individual who can assign meanings to outputs [24]. This means that the person interacting with AI can understand its functionality, purpose, or impact on the context of use. Therefore, the process of interpretation involves mapping an abstract concept, such as a predicted class or category, into a domain that is within the grasp of human understanding [58].
@@ 3.1.2 Fairness
In AI systems, Fairness reflects the concepts of equality and inclusiveness to avoid biases and discriminatory behaviors safeguarding fundamental human rights and values [24]. An approach that considers humans as a whole is essential for ensuring that AI can contribute positively to society and protect them against potential harms [2]. Fairness is characterized by two dimensions: Rightful Information and Non-Discrimination.
@@
AI systems must disseminate accurate and reliable information to minimise the risk of incorrect or incomplete knowledge and to mitigate the risk of manipulation and persuasion of humans [22, 68]. Providing accurate and complete information is essential to explain AI decisions to safeguard individual rights and freedom, foster a sense of trust and understanding among humans, and enhance the ethical use of AI systems [79].
@@
Although AI is commonly used to boost productivity through automation, it is important to ensure that models are trained in accordance with ethical and societal norms that minimize discriminatory behaviors, avoiding that individuals are treated differently or unequally without any justifiable reason [5]. Thus, the whole pipeline of creating AI-based systems must be monitored and checked since biases can rise unintentionally during the early training phases of models [51, 58].
@@ 3.1.3 Automation Level
As AI becomes increasingly integrated into countless aspects of human life, studying the appropriate balance between automation and human control in human-AI interactions is necessary [12]. Although there are contexts of use in which humans need or wish for fully automated systems in which their control is not necessary, it is important to address the ethical and legal consequences of undesired events caused by AI systems’ outputs. This implies that automation can be considered as a spectrum and not as a binary feature [13]. Automation Level is characterized by two dimensions: Human-on-the-loop and Human-in-the-loop.
@@
Providing humans with appropriate oversight when interacting with AI systems is necessary to enable them to check, monitor, and supervise the system’s behaviour. Oversight is a precondition to allow human intervention, which guarantees AI-assisted decisions rather than AI-driven decisions. This can avoid irreversible consequences and minimise risk and biased outputs while safeguarding human rights [9]. Human oversight is strictly related to Transparency, as AI models must provide effective explanations to users, ensuring that users can effectively interpret outputs in order to properly modify their behaviour, if necessary [28].
@@
Designing AI systems emphasising human control is useful in situations where humans need to actively participate in the decision-making process. In this context, the concept of Controllable AI is introduced, which reinforces the importance of human control to detect malfunctions and recover from dangerous situations; controlling the behavior of AI means influencing its output and processes in accordance with the context of use and human expertise for a more safe and reliable interaction [31]. Some of the conditions must be in place for humans to take control of the interaction properly: appropriate algorithmic transparency of AI models and high levels of feedback and affordance in User Interfaces are needed since communication is key in any kind of relationship [67].
@@ 3.1.4 Protection
The human-centric approach undertaken by the AI Act aims at ensuring that users are safeguarded against harm, threats, or intrusion. This principle is strongly intertwined with the legal requirements set by governmental norms and rules that designers, developers, and deployers must comply with to protect users from unsafe behavior. There is a need to create secure and resilient AI systems that can preserve users’ privacy. In this regard, the AI Act recalls the General Data Protection Regulation (GDPR), which emphasized the integration of privacy and data protection into the design and development of systems [14]. Protection refers to a large spectrum of aspects, laying its ground in cybersecurity, and being characterized by three dimensions: Privacy, Safety, and Security.
@@
An AI system that fosters users’ privacy can safeguard individuals’ sensitive data from improper access, theft, or loss [15]. This principle transcends from the specific case of AI systems, as applying the proper techniques to preserve data and protect individuals’ identities impacts multiple system components and requires developers to inform end-users about how their data is being handled [58]. The correct application of data preservation techniques and protective measures can have far-reaching impacts on multiple system components.
@@
A safe system is designed to fulfill its intended function without causing harm to living beings or the environment [15]. This concept is linked to the well-being and welfare of humans affected by AI and is also connected to the system’s level of automation. The objective is to mitigate risk and prevent accidents by removing barriers to error reporting and fostering a collaborative and communicative environment. This ensures that end-users are always informed about potentially harmful practices that could threaten their rights [50].
@@
To protect individuals, systems must be secure by design, implying that they must incorporate solutions that allow management, monitoring, and recovery from external threats [51]. This dimension is relevant to all types of systems, regardless of AI features, since security implies the preservation of the CIA triad (Confidentiality, Integrity, and Availability), meaning that strong prevention and recovery measures must be implemented. AI systems, especially those categorized as high-risk by the AI Act, should be designed to be resilient against attacks and consistently perform securely throughout their lifecycle since they deal with extremely private and sensitive information about their users [67].
3.2 Properties of Human-Centred AI
Three properties emerged from the SLR, which underlie the principles that were identified and that must characterise AI-based systems that are designed and developed compliantly with the AI Act. The analysis was performed using hierarchical clustering, and after the initial step of identifying the first set of clusters, a hierarchical grouping was brought to the identification of the principles described in the previous section. A further grouping step allowed us to identify the overall properties of a truly human-AI symbiotic relationship: Trustworthiness, Robustness, and Sustainability. This section details these properties, highlighting their relationship with the four principles described in Section 3. It is highlighted that these properties are not a sufficient condition for AI systems to be symbiotic, but they emerged as crucial components of the design phase from this SLR, influenced by its goals. Figure 3 shows the connection among the identified properties and principles.
@@ 3.2.1 Trustworthiness
Trustworthiness is a critical, frequently-discussed topic in the context of the human-centric approach of the AI Act. It is essential for high-risk systems because it is necessary for reliable, safe, and positive interactions [70]. It can be fostered by implementing appropriate transparency techniques in AI models in order to make users aware of the motivations that lie behind outputs and responses [35]. This property is still a subject of debate within the research community and the legal landscape. For example, in the context of HCAI, a trustworthy system is one that deserves human trust, implying that it must align with users’ needs, preferences, and cognitive models to achieve successful interactions [66]. On the other hand, the AI High Level Expert Group (HLEG) identifies trustworthiness as the umbrella property to ensure a human-centric approach to AI [23]. Nevertheless, our analysis showed that trustworthiness is an important property, which covers the four principles, but it is not a sufficient condition for the establishment of a symbiotic relationship [11]. It emerges that an AI system is trustworthy if it enables humans to properly oversee and/or control its performance with an appropriate automation level while exhibiting fair behavior and respecting humans in all their dimensions.
@@ 3.2.2 Robustness
Referring to the AI Act, the Robustness of AI has been defined as their ability to perform reliably and effectively under various conditions, including unexpected or challenging ones. Robustness is mentioned in Article 15 of the law, which is titled “Accuracy, Robustness and Cybersecurity” in the context of ensuring that high-risk systems feature fail-safe plans and technical redundancy solutions [14]. The motivations behind this lie in the fact that a robust AI system exhibits a safe, resilient, and reliable behaviour, which fosters trust in humans, thus being one of the factors that contribute to a symbiotic human-centred relationship [35, 73].
@@ 3.2.3 Sustainability
Sustainability is a factor that deployers should consider when creating systems to align with the EU’s goals for the near future. As environmental concerns are substantially increasing worldwide and given the EU’s efforts in trying to reduce emissions and energy consumption, a sustainable approach is necessary to minimize the environmental impact of AI and to create long-lasting products that uphold ethical standards [55, 67]. For example, reducing the complexity of Neural Networks can benefit AI systems, as it reduces the required computational power in the training phase while increasing transparency [63]. Although there are best practices that designers and developers should follow and requirements set by the law, it is highlighted that there is a lack of standardized methodologies for evaluating the impact of AI systems in terms of sustainability [19]. As the AI Act was refined, the EU has made steps forward in this direction by releasing compliance-checkers and user-friendly explorers of the law, but they remain auto-assessment tools that do not foster compliance by design. AI systems must comply with the law and satisfy the requirements in terms of sustainability, being reliable and resilient regardless of the environmental and contextual changes that they undergo.

### v3

@@ 1 Introduction
This work aims to identify the principles that can guide the design of SAI systems to comply with the current legal landscape considering humans as the core of the development process [30]. With this goal in mind, we performed a Systematic Literature Review (SLR), employing content analysis to understand what researchers suggest about the relationship between SAI and the AI Act. The main contribution of this study is the proposal of principles to guide stakeholders in the design of SAI systems, i.e., of HCAI systems that favour a symbiotic relationship with humans. It is finally put in the context of the relevant literature, comparing it with other principled-based approaches to SAI and HCAI. The goal of this study is thus better formulated in the following research question:
@@
The article is structured as follows: Section 2 describes the methodology adopted to conduct the SLR; Section 3 presents the results obtained by the SLR (i.e., principles and properties that characterize the human-AI symbiosis); Section 4 discusses the analysis of the literature presenting the challenges, trends, and limitations faced while conducting the SLR.
2 Methodology
@@
    european AND (“artificial intelligence” OR ai) AND act AND (human OR human-centric OR human-centred OR human-centered)
@@ 2.2 Coding and Classification
The papers were classified using a mixed approach of data coding: i.e., a-priori and in-vivo. A-priori coding consists of the categorisation of papers with codes that are established before the classification process. In contrast, in-vivo coding involves the definition of the codes as the papers are read and analysed [37]. First, articles were classified according to the a-priori defined codes to guarantee coherence in the analysis. Then, in-vivo coding was adopted to make the methodology feasible and integrate new and emerging concepts. During this phase, the name of each code, along with its sub-codes, was carefully and iteratively refined with respect to the topics that characterize the specific group of articles.
The name of each principle and its dimensions correspond to the name of codes and sub-codes defined during the process. They were collectively chosen by the researchers who took part in the SLR, reflecting terms that are recurrent in the set of papers and those used in the regulation.
@@ 2.3 Conducting the SLR
The first step of the execution was searching for academic articles about the topic by running the five queries on the Scopus digital library because it is comprehensive and includes the most relevant and accredited digital libraries, including journal and conference manuscripts. Each query provided the following results: Q1: 530, Q2: 223, Q3: 4, Q4: 1, and Q5: 4. It emerged that Q1 and Q2 returned a higher number of papers obtained since they are broader and more exploratory than the others.
@@
From a total of 762 papers, after a check for duplicates (i.e., 158 files), 604 publications were obtained. A more in-depth step was taken, analysing titles and keywords, which resulted in the removal of 365 papers since they did not match the inclusion criteria. In the end, each publication was further analysed by reviewing the abstract, the introduction, and the conclusions, obtaining the final set of papers. The final set used for the literature review contains 58 papers. Figure 1 shows the PRISMA diagram of the whole process [59].
@@
3 Results: A Principled Human-Centred AI Framework to Build SAI
The four principles, represented in Figure 2, and their relationship with the additional three properties are detailed in the following sections, discussing how they can support the creation of SAI systems, i.e., HCAI systems that foster a symbiotic relationship with humans [12].
The principles emerged during the coding phase. In a first iteration of the SLR, the report by [Fjeld2020Principled] was taken as the a-priori coding scheme since it presents a principle-based framework for AI systems before the release of the AI Act. In order to determine how the regulation impacted the creation of such systems, the coding scheme was modified during later iterations through the in-vivo technique. After the analysis of the articles and the definition of the final codes, additional intrinsic and transversal aspects emerged, which were identified as properties.
@@ 3.1 Principles for Human-AI Symbiosis
Defining principles to guide the design and development of AI systems through a human-centred approach is crucial to establish a symbiotic relationship between the two parties [Desolda2024Humancentered]. It is important to guarantee that AI systems adhere to the highest standards of ethical conduct. Thus, designers and developers must ensure that the implemented solutions comply with laws and regulations safeguarding human rights and European values, making these systems safer for end-users [Helberger2023ChatGPT, Worsdorfer2023Mitigating]. To this end, providers must register high-risk AI models in an EU database managed by the European Commission to enhance public transparency and enable oversight by authorities [Pavlidis2024Unlocking].
@@ 3.1.1 Transparency
Transparency can guarantee that AI systems are effectively overseen by humans and allow intervention when potential harm occurs [Pavlidis2024Unlocking]. Transparency ensures that critical information about how the AI model was trained and structured is available to humans [Pavlidis2024Unlocking], since stakeholders must understand how AI models function and the reasoning behind their decisions to be able to intervene [Pavlidis2024Unlocking]. Being a multi-faceted property that concerns AI models, their components, and algorithms, Transparency serves as an umbrella term to the functional understanding of the model and the rationale behind its operations [Gyevnar2023Bridging, Urquhart2022Legal]. It is characterized by two dimensions: Explainability and Interpretability [HighLevelExpertGrouponArtificialIntelligence2020Assessment].
@@
Explainability aims to provide explanations concerning AI systems’ operations, ensuring that even when humans cannot understand how an AI system provides an output, they can at least receive information about why it was produced [Siegel2024Media]. In other words, this aspect contributes to the degree to which an AI system is open and observable to humans [Pavlidis2024Unlocking, Urquhart2022Legal]. When interacting with AI systems, humans should have the right to explanations to make informed decisions, supported by AI and not replaced by it [Malgieri2024Licensing, Wagner2024Navigating], while being able to oversee its processes [Tartaro2023European].
@@
An AI system can be considered interpretable if it can be correctly understood by an individual who can assign meanings to outputs [HighLevelExpertGrouponArtificialIntelligence2020Assessment]. This means that the person interacting with AI can understand its functionality, purpose, or impact on the context of use. Therefore, the process of interpretation involves mapping an abstract concept, such as a predicted class or category, into a domain that is within the grasp of human understanding [Pavlidis2024Unlocking].
@@ 3.1.2 Fairness
In AI systems, Fairness reflects the concepts of equality and inclusiveness to avoid biases and discriminatory behaviors safeguarding fundamental human rights and values [HighLevelExpertGrouponArtificialIntelligence2020Assessment]. An approach that considers humans as a whole is essential for ensuring that AI can contribute positively to society and protect them against potential harms [Anamaria2023Artificial]. Fairness is characterized by two dimensions: Rightful Information and Non-Discrimination.
@@
AI systems must disseminate accurate and reliable information to minimise the risk of incorrect or incomplete knowledge and to mitigate the risk of manipulation and persuasion of humans [Herbosch2024Fraud, Sposini2024Neuromarketing]. Providing accurate and complete information is essential to explain AI decisions to safeguard individual rights and freedom, foster a sense of trust and understanding among humans, and enhance the ethical use of AI systems [Worsdorfer2023Mitigating].
@@
Although AI is commonly used to boost productivity through automation, it is important to ensure that models are trained in accordance with ethical and societal norms that minimize discriminatory behaviors, avoiding that individuals are treated differently or unequally without any justifiable reason [Biewer2024Software]. Thus, the whole pipeline of creating AI-based systems must be monitored and checked since biases can rise unintentionally during the early training phases of models [Nikiforov2024Groups, Pavlidis2024Unlocking].
@@ 3.1.3 Automation Level
As AI becomes increasingly integrated into countless aspects of human life, studying the appropriate balance between automation and human control in human-AI interactions is necessary [Enqvist2023Human]. Although there are contexts of use in which humans need or wish for fully automated systems in which their control is not necessary, it is important to address the ethical and legal consequences of undesired events caused by AI systems’ outputs. This implies that automation can be considered as a spectrum and not as a binary feature [Esposito2024Fine]. Automation Level is characterized by two dimensions: Human-on-the-loop and Human-in-the-loop.
@@
Providing humans with appropriate oversight when interacting with AI systems is necessary to enable them to check, monitor, and supervise the system’s behaviour. Oversight is a precondition to allow human intervention, which guarantees AI-assisted decisions rather than AI-driven decisions. This can avoid irreversible consequences and minimise risk and biased outputs while safeguarding human rights [CoveloDeAbreu2024Artificial]. Human oversight is strictly related to Transparency, as AI models must provide effective explanations to users, ensuring that users can effectively interpret outputs in order to properly modify their behaviour, if necessary [Hupont2022Landscape].
@@
Designing AI systems emphasising human control is useful in situations where humans need to actively participate in the decision-making process. In this context, the concept of Controllable AI is introduced, which reinforces the importance of human control to detect malfunctions and recover from dangerous situations; controlling the behavior of AI means influencing its output and processes in accordance with the context of use and human expertise for a more safe and reliable interaction [Kieseberg2023Controllable]. Some of the conditions must be in place for humans to take control of the interaction properly: appropriate algorithmic transparency of AI models and high levels of feedback and affordance in User Interfaces are needed since communication is key in any kind of relationship [Siegel2024Media].
@@ 3.1.4 Protection
The human-centric approach undertaken by the AI Act aims at ensuring that users are safeguarded against harm, threats, or intrusion. This principle is strongly intertwined with the legal requirements set by governmental norms and rules that designers, developers, and deployers must comply with to protect users from unsafe behavior. There is a need to create secure and resilient AI systems that can preserve users’ privacy. In this regard, the AI Act recalls the General Data Protection Regulation (GDPR), which emphasized the integration of privacy and data protection into the design and development of systems [EuropeanParliament2024Regulation]. Protection refers to a large spectrum of aspects, laying its ground in cybersecurity, and being characterized by three dimensions: Privacy, Safety, and Security.
@@
An AI system that fosters users’ privacy can safeguard individuals’ sensitive data from improper access, theft, or loss [Fjeld2020Principled]. This principle transcends from the specific case of AI systems, as applying the proper techniques to preserve data and protect individuals’ identities impacts multiple system components and requires developers to inform end-users about how their data is being handled [Pavlidis2024Unlocking]. The correct application of data preservation techniques and protective measures can have far-reaching impacts on multiple system components.
@@
A safe system is designed to fulfill its intended function without causing harm to living beings or the environment [Fjeld2020Principled]. This concept is linked to the well-being and welfare of humans affected by AI and is also connected to the system’s level of automation. The objective is to mitigate risk and prevent accidents by removing barriers to error reporting and fostering a collaborative and communicative environment. This ensures that end-users are always informed about potentially harmful practices that could threaten their rights [Neuwirth2023Prohibited].
@@
To protect individuals, systems must be secure by design, implying that they must incorporate solutions that allow management, monitoring, and recovery from external threats [Nikiforov2024Groups]. This dimension is relevant to all types of systems, regardless of AI features, since security implies the preservation of the CIA triad (Confidentiality, Integrity, and Availability), meaning that strong prevention and recovery measures must be implemented. AI systems, especially those categorized as high-risk by the AI Act, should be designed to be resilient against attacks and consistently perform securely throughout their lifecycle since they deal with extremely private and sensitive information about their users [Siegel2024Media].
3.2 Properties of Human-AI Symbiosis
Three properties emerged from the SLR, which underlie the principles that were identified and that must characterise AI-based systems that are designed and developed compliantly with the AI Act. After the coding phase, an additional meta-analysis was performed using hierarchical clustering, and after the initial step of identifying the first set of clusters, a hierarchical grouping was brought to the identification of the principles described in the previous section. A further grouping step allowed us to identify the overall properties of a truly human-AI symbiotic relationship: Trustworthiness, Robustness, and Sustainability. This section details these properties, highlighting their relationship with the four principles described in Section 3. It is highlighted that these properties are not a sufficient condition for AI systems to be symbiotic, but they emerged as crucial components of the design phase from this SLR, influenced by its goals. Figure 3 shows the connection among the identified properties and principles.
@@ 3.2.1 Trustworthiness
Trustworthiness is a critical, frequently-discussed topic in the context of the human-centric approach of the AI Act. It is essential for high-risk systems because it is necessary for reliable, safe, and positive interactions [Stettinger2024Trustworthiness]. It can be fostered by implementing appropriate transparency techniques in AI models in order to make users aware of the motivations that lie behind outputs and responses [Laux2024Trustworthy]. This property is still a subject of debate within the research community and the legal landscape. For example, in the context of HCAI, a trustworthy system is one that deserves human trust, implying that it must align with users’ needs, preferences, and cognitive models to achieve successful interactions [Shneiderman2022HumanCentered]. On the other hand, the AI High Level Expert Group (HLEG) identifies trustworthiness as the umbrella property to ensure a human-centric approach to AI [High-LevelExpertGrouponArtificialIntelligence2019Ethics]. Nevertheless, our analysis showed that trustworthiness is an important property, which covers the four principles, but it is not a sufficient condition for the establishment of a symbiotic relationship [Desolda2024Humancentered]. It emerges that an AI system is trustworthy if it enables humans to properly oversee and/or control its performance with an appropriate automation level while exhibiting fair behavior and respecting humans in all their dimensions.
@@ 3.2.2 Robustness
Referring to the AI Act, the Robustness of AI has been defined as their ability to perform reliably and effectively under various conditions, including unexpected or challenging ones. Robustness is mentioned in Article 15 of the law, which is titled “Accuracy, Robustness and Cybersecurity” in the context of ensuring that high-risk systems feature fail-safe plans and technical redundancy solutions [EuropeanParliament2024Regulation]. The motivations behind this lie in the fact that a robust AI system exhibits a safe, resilient, and reliable behaviour, which fosters trust in humans, thus being one of the factors that contribute to a symbiotic human-centred relationship [Laux2024Trustworthy, Tallberg2024AI].
@@ 3.2.3 Sustainability
Sustainability is a factor that deployers should consider when creating systems to align with the EU’s goals for the near future. As environmental concerns are substantially increasing worldwide and given the EU’s efforts in trying to reduce emissions and energy consumption, a sustainable approach is necessary to minimize the environmental impact of AI and to create long-lasting products that uphold ethical standards [Pagallo2022Environmental, Siegel2024Media]. For example, reducing the complexity of Neural Networks can benefit AI systems, as it reduces the required computational power in the training phase while increasing transparency [Schwartz2020Green]. Although there are best practices that designers and developers should follow and requirements set by the law, it is highlighted that there is a lack of standardized methodologies for evaluating the impact of AI systems in terms of sustainability [Hacker2023European]. As the AI Act was refined, the EU has made steps forward in this direction by releasing compliance-checkers and user-friendly explorers of the law, but they remain auto-assessment tools that do not foster compliance by design. AI systems must comply with the law and satisfy the requirements in terms of sustainability, being reliable and resilient regardless of the environmental and contextual changes that they undergo.
In summary, through content analysis performed on 58 articles, four principles were identified as requirements for designing AI Act-compliant SAI systems: Transparency, Fairness, Automation Level, and Protection. Through further analysis, three additional properties were identified: Trustworthiness, Robustness, and Sustainability.